OpenStudy (anonymous):

find the zeros of each function. state the multiplicity of multiple zeros y=(x-3)^2(x+1)

OpenStudy (aum):

y = (x-3)^2 * (x+1) = 0 (x-3)^2 = 0 OR (x+1) = 0 x = 3 OR x = -1 multiplicity = 2 multiplicity = 1

OpenStudy (anonymous):

thank you

OpenStudy (aum):

you are welcome.

OpenStudy (anonymous):

multiplicity of y=x^2+3x+2